Question-In the equation ax^2 + bx + c=0, the value of b^2 -4ac is called the of the quadratic equation. What does the value tell you about the real roots of the equation?
Part 2, they don't want the graph. Given the equation,determine how many x intercepts the parabola has. And whether its vertix lies above or below or on the axis.
Thanks appreciate all your help. Could not do it without you.

Answer by stanbon(75887)

In the equation ax^2 + bx + c=0, the value of b^2 -4ac is called the determinant of the quadratic equation.
What does the value tell you about the real roots of the equation?
Ans:
If b^2-4ac < 0 there are no Real Number roots.
If b^2-4ac = 0 there are two equal Real Number roots.
If b^2-4ac > 0 there are two unequal Real Number roots.
--------------------------------------------------------
Part 2, they don't want the graph. Given the equation,determine how many x intercepts the parabola has. And whether its vertix lies above or below or on the axis.
If b^2-4ac < 0 there are no x-intercepts; vertex above or below and graph
does not intersect the x-axis.
If b^2-4ac = 0 the graph touches the x-axis at one point: vertex on x-axis
If b^2-4ac > 0 the graph intersects the x-axis at two points: vertex above or below and graph passes thru the x-axis.
=================================
Cheers,
Stan H.